OpenStudy (dls):

The potential energy between two toms(in a molecule) is given by.. U(x)=a/x^12-b/x^6.. where a & b=+ve and x=distance b/w 2 atoms..The molecule is in stable equilibrium when x=?

OpenStudy (dls):

\[\LARGE (ax^{-12} -bx^{-6})\] 2nd derivative=minimum \[\LARGE \frac{156}{x^{14}}-\frac{42b}{x^8}<0\] if im going right

OpenStudy (dls):

2nd derivative=negative*

OpenStudy (shubhamsrg):

whats the need? body will be in eqbm when net force on it = 0

OpenStudy (shubhamsrg):

F= -dU/dx

OpenStudy (dls):

yes

OpenStudy (shubhamsrg):

So just make the substituion.

OpenStudy (dls):

Diff U?

OpenStudy (dls):

once?

OpenStudy (shubhamsrg):

no, diff yourself! >.<

OpenStudy (shubhamsrg):

yes once

OpenStudy (dls):

\[\LARGE -12ax^{-13}+6bx^{-7}\]

OpenStudy (dls):

=F

OpenStudy (shubhamsrg):

=-F

OpenStudy (shubhamsrg):

it doesn;t matter, just put that =0

OpenStudy (dls):

0?

OpenStudy (shubhamsrg):

net force =0

OpenStudy (dls):

that is neutral eqm?

OpenStudy (dls):

net force=-ve is stable eqm

OpenStudy (shubhamsrg):

hein ?

OpenStudy (shubhamsrg):

just put that 0, >.<

OpenStudy (dls):

\[\LARGE x=\sqrt[7]{\frac{2a}{b}} \] nice

OpenStudy (dls):

6*

OpenStudy (shubhamsrg):

:)

OpenStudy (shubhamsrg):

now lets dance? :)

OpenStudy (dls):

but why 0

OpenStudy (shubhamsrg):

celebration time!

OpenStudy (shubhamsrg):

I can sing also, I know tinku jiya song :)

OpenStudy (dls):

.-.

OpenStudy (shubhamsrg):

pal pal na maane.. ??

OpenStudy (dls):

go back where you came from right now :P

OpenStudy (shubhamsrg):

my space ship left me! :( I am right now staying with rohit.

OpenStudy (shubhamsrg):

please complete the song, pal pal na maane.. ?

OpenStudy (dls):

\[\LARGE \rho hit ?\]

OpenStudy (shubhamsrg):

phuck ?

OpenStudy (dls):

D:

OpenStudy (shubhamsrg):

pal pal na maane??? >.<

OpenStudy (dls):

?

OpenStudy (dls):

lolwut

OpenStudy (dls):

gtfo.ur job is done :D

OpenStudy (shubhamsrg):

http://meme.andreavenditti.com/okay.png

OpenStudy (shubhamsrg):